OpenStudy (anonymous):

how would i solve 48=4x+2y 48=x+y

OpenStudy (anonymous):

divide the first equation by by 2 to simplify then subtract to eliminate x first 24=2x+y -48= x+y -------- -24=x hence 48=-24+y 48+24=y 72=y hence x=-24 y=72

OpenStudy (anonymous):

what is hence?

OpenStudy (anonymous):

x=36 and y=60